# Partition functions of determinantal point processes on polarized Kähler manifolds

### v3

@@ 3. Asymptotic expansion of the DPP partition function
$\tilde{S}_{\phi}[\varphi_{1},\varphi_{0}]:=\int_{M}BC(\phi;\omega_{1},\omega_{0})\frac{\omega_{0}^{n+1-j}}{(n+1-j)!}+\phi(R_{1})\frac{-i}{(n+1-j)!}\sum^{n-j}_{s=0}(\varphi_{1}-\varphi_{2})\omega_{1}^{s}\wedge\omega_{0}^{n-j-s}$ (3.22)
satisfies cocycle identity without polarization assumption on $[\omega]$. Functionals of similar type were considered in [Don85, Don87, Tia00a]. Note that $\tilde{S}_{\phi}$ can be easily modified to become a functional on the space of Kähler forms, as we now demonstrate.
@@
This can be easily verified from either (3.8) or (3.10). We have an alternative expression for the asymptotics (3.11) where each of the coefficients is given by a geometric functional in the sense that it only depends on Kähler forms. Let $V:=\int_{M}\omega^{n}/n!$ be a volume of $(M,\omega)$.
Theorem 3.2 (Asymptotics of the Partition Functions, Version II).
$(2\pi)^{n}\log\frac{Z_{k}[\varphi]}{Z_{k}[0]}=kd_{k}(2\pi)^{n}S_{0}[\varphi,0]+k^{n}S_{1}[\omega_{\varphi},\omega]+k^{n-1}S_{2}[\omega_{\varphi},\omega]+k^{n-2}S_{3}[\omega_{\varphi},\omega]+\cdots.$ (3.24)
  1. (1)
    $S_{j}[\omega_{1},\omega_{0}]$ $=-S_{j}[\omega_{0},\omega_{1}],$ (3.25)
    $S_{j}[\omega_{2},\omega_{0}]$ $=S_{j}[\omega_{2},\omega_{1}]+S_{j}[\omega_{1},\omega_{0}];$ (3.26)
  2. (2)
    $\widehat{a_{j}(\omega_{\varphi})}:=\frac{1}{V}\int_{M}a_{j}(\omega_{\varphi})\frac{\omega_{\varphi}^{n}}{n!}=\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L),$ (3.28)
    which does not depend on $\omega_{\varphi}$.
  3. (3)
    $S_{j}[\omega_{2},\omega_{1}]=s_{j}[\omega_{2}]-s_{j}[\omega_{1}],$ (3.29)
Proof.
@@
$(2\pi)^{n}\log\frac{Z_{k}[\varphi]}{Z_{k}[0]}=kd_{k}(2\pi)^{n}S_{0}[\varphi,0]+k^{n}S_{1}[\varphi,0]+k^{n-1}S_{2}[\varphi,0]+\cdots,$ (3.30)
@@
$S_{0}[\varphi,0]:=\frac{1}{V}\widetilde{S_{0}}[\varphi,0]$ (3.31)
@@
$S_{j}[\varphi+c,0]=\widetilde{S_{j}}[\varphi+c,0]-\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L)\times\widetilde{S_{0}}[\varphi+c,0]$ (3.33)
@@
$\qquad-\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L)\times\widetilde{S_{0}}[\varphi,0]+c\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L)$ (3.34)
$=S_{j}[\varphi,0]$ (3.35)
for $j>0$. Thus, the functionals $S_{j}[\cdot,0],\,j>0$ on $\mathcal{K}_{\omega}$ descend to the fucntionals on Kähler metrics $\mathcal{K}_{0}=\mathcal{K}_{\omega}/\mathbb{R}$. Cocyle identity and exact cocycle property for $j>n+1$ follow from corresponding properties of $\widetilde{S_{j}}$, and (2) follows from (3.32). ∎
From (2.4) and (3.27), one can see that $S_{0}$ is (minus of) the Aubin-Yau functional $I$, and $S_{1}$ is the Mabuchi functional $\mathcal{M}$ in Kähler geometry. It can also be directly checked by comparing the explicit formula for $I,\mathcal{M}$ [PS07] with (3.10) and (3.32). For $j=2$, since $\beta_{1}$ only depends on $[\omega]$, $S_{2}$ is explicitly given by
@@
$+\frac{i}{V}\left(\int_{M}\operatorname{Td}_{2}(T^{1,0}M)\frac{\omega^{n-2}}{(n-2)!}\right)\times\int_{M}\frac{-i}{(n+1)!}\sum^{n}_{s=0}\varphi\omega_{\varphi}^{s}\wedge\omega^{n-s}.$ (3.36)
@@
$\delta_{\varphi}S_{2}[\omega_{\varphi},\omega]=\frac{1}{6}\int_{M}\delta\varphi\Delta_{\varphi}R_{\varphi}\,\omega_{\varphi},$ (3.38)
which is the variation of the Liouville action restricted to fixed area metrics, written in terms of Kähler gauge [Kle14, (5.10)]. Therefore, $S_{2}$ naturally generalizes the Liouville action to all dimensions. We refer to $S_{2}$ as the generalized Liouville action, and we will compute its second variation in Appendix A; see Proposition A.1.
@@
In [Kle14, (6.6)], Klevtsov explicitly computed the asysmptotics of $\log\frac{Z_{k}[\varphi]}{Z_{k}[0]}$ up to $j\leq 4$ in dimension $n=1$. Based on that, it was conjectured that for the general dimension $n$, the first $n+2$ coefficients are nontrivial action functionals (satisfying cocycle identity) and terms with $j>n+1$ are exact cocycles. Thus, Theorem 3.2 confirms this conjecture. It was also asked what is the relationship between $S_{2}$ and known functionals in Kähler geometry, such as Chen-Tian functional $E_{1}$ [CT02]. In fact, from (3.37), one can see that $S_{2}$ is a linear combination of the Chen-Tian functional $E_{1}$ and the Bando-Mabuchi functional $\mathcal{M}_{2}$ [BM86]. By the results of [BM86, CT02], one can observe that $S_{2}$ integrates a Futaki-type holomorphic invariant. We will generalize this for all $j>0$ in the next section.
Note that (3.8) and (3.10) make sense as formal expressions for a nonpolarized Kähler class $[\omega]$. We expect that they are equal in general, without a polarization assumption. In particular, it would imply that the 1-forms defined on $\mathcal{K}_{\omega}$ by
@@
for $\psi\in T_{\varphi}\mathcal{K}_{\omega}=C^{\infty}(M,\mathbb{R})$, are closed. It is known that this holds for $j=0,1$ by [Mab86], and for $j=2$ by [BM86, CT02]. We will verify these claims for $j=2$ in Appendix A by direct computation.
@@ 4. Holomorphic invariants from Bergman kernel asymptotics and critical points
In this section, we show that there is an obstruction to the existence of the critical points for each $S_{j}$, given by a holomorphic invariant introduced by Futaki [Fut04]. Denote by $\mathfrak{h}(M)$ the Lie algebra of holomorphic vector fields on $(M,\omega)$. For $X\in\mathfrak{h}(M)$, by Hodge theory, there is a unique harmonic $(0,1)$-form $\tau$ and function $\theta_{X}$ satisfying
@@
Then $\theta_{X}$ is defined up to an additive constant, and $\tau$ can be assumed to be zero without loss of generality for our purposes. Consider $\nabla X=X^{q}_{p}\frac{\partial}{\partial z^{q}}\otimes dz^{p}$ as an $\operatorname{End}(T^{1,0}M)$-valued 00-form on $M$, where
$X^{q}_{p}=\frac{\partial X^{q}}{\partial z^{p}}+\Gamma^{q}_{pr}X^{r}.$ (4.2)
We will need the following formula.
@@
$\iota_{X}R=-\bar{\partial}\nabla X,$ (4.3)
where $R$ is the curvature 2-form of $\omega$.
@@
¯$\bar{\partial}\nabla X$ =¯$=\bar{\partial}\left(X^{q}_{p}\right)=\frac{\partial X^{q}_{p}}{\partial\bar{z}^{l}}d\bar{z}^{l}=\left(\frac{\partial^{2}X^{q}}{\partial z^{p}\partial\bar{z}^{l}}+\frac{\partial\Gamma^{q}_{pr}}{\partial\bar{z}^{l}}X^{r}+\Gamma^{q}_{pr}\frac{\partial X^{r}}{\partial\bar{z}^{l}}\right)d\bar{z}^{l}$ (4.4)
$=\frac{\partial\Gamma^{q}_{pr}}{\partial\bar{z}^{l}}X^{r}d\bar{z}^{l}=-R_{p}{}^{q}{}_{r\bar{l}}X^{r}d\bar{z}^{l}=-\iota_{X}R.$ (4.5)
Note that by taking $\operatorname{Tr}_{\operatorname{End}(T^{1,0}M)}$ of (4.3), we get
$\iota_{X}\operatorname{Ric}=\bar{\partial}\Delta\theta_{X}.$ (4.6)
Now we prove the main result in this section. It generalizes formula [Lu04, (4.4)] for all $j\geq 0$.
@@
For non-purely imaginary $\theta_{X}$, (4.7) holds with $\theta_{X}$ replaced by $i\,{\rm Im\,}\theta_{X}$ and $\nabla X$ replaced by ${\rm Re\,}\nabla X$.
@@
Assume $\theta_{X}$ to be purely imaginary. Let $f_{t}\in Aut(M)$ be the flow of ${\rm Re\,}X=(X+\overline{X})/2$. Let $\omega_{t}:=f_{t}^{*}\omega$. Then $\omega_{t}$ is given by $\omega_{t}=\omega+i\partial\bar{\partial}\varphi_{t}$, with $\dot{\varphi_{t}}=if^{*}_{t}\theta_{X}$; see [Wei02, p8].
Consider $\left.-\frac{d}{dt}\right|_{t=0}\widetilde{S_{j}}[\omega_{t},\omega]$. From the first variation formula (3.14), we have
$-\frac{d}{dt}$ $\widetilde{S_{j}}[\omega_{t},\omega]=\int_{M}\dot{\varphi_{t}}\left(a_{j}(\omega_{t})-\Delta_{t}a_{j-1}(\omega_{t})\right)\frac{\omega_{t}^{n}}{n!}=\int_{M}if_{t}^{*}\theta_{X}\left(a_{j}(\omega_{t})-\Delta_{t}a_{j-1}(\omega_{t})\right)\frac{\omega_{t}^{n}}{n!}$
$=i\int_{M}f_{t}^{*}\theta_{X}f_{t}^{*}\left(a_{j}(\omega)-\Delta a_{j-1}(\omega)\right)\frac{f_{t}^{*}\omega^{n}}{n!}=i\int_{M}\theta_{X}\left(a_{j}(\omega)-\Delta a_{j-1}(\omega)\right)\frac{\omega^{n}}{n!}$ (4.8)
for all $t\in\mathbb{R}$.
Alternatively, from (3.10) we have
$\left.-\frac{d}{dt}\right|_{t=0}\widetilde{S_{j}}[\omega_{t},\omega]$ $=i\int_{M}\left.\frac{d}{dt}\right|_{t=0}BC(\operatorname{Td}_{j};\omega_{t},\omega)\frac{\omega^{n+1-j}}{(n+1-j)!}$
$\qquad+\operatorname{Td}_{j}(R)\left.\frac{d}{dt}\right|_{t=0}\frac{-i}{(n+1-j)!}\sum^{n-j}_{s=0}\varphi_{t}\omega_{t}^{s}\wedge\omega^{n-j-s},$ (4.9)
Now we compute $\frac{d}{dt}BC(\cdot;\omega_{t},\omega)$ using property (3) of Bott-Chern forms (Proposition 2.2). First, note that
$\left.\dot{\omega_{t}}\right|_{t=0}=\left.i\partial\bar{\partial}\dot{\varphi_{t}}\right|_{t=0}=\left.i\partial\bar{\partial}if_{t}^{*}\theta_{X}\right|_{t=0}=-\partial\bar{\partial}\theta_{X}=\partial(\iota_{X}\omega).$ (4.10)
This implies, by local computation,
$\left.\dot{\omega_{t}}{\omega_{t}}^{-1}\right|_{t=0}=\frac{\partial}{\partial z^{p}}\left(X^{s}g_{s\bar{k}}\right)g^{q\bar{k}}=\frac{\partial X^{s}}{\partial z^{p}}\delta^{q}_{s}+X^{s}\frac{\partial g_{s\bar{k}}}{\partial z^{p}}g^{q\bar{k}}=\frac{\partial X^{q}}{\partial z^{p}}+X^{s}\Gamma^{q}_{sp}=X^{q}_{p}=\nabla X.$ (4.11)
Thus we have
Combining (4.12) and (4.13), we obtain
$\left.-\frac{d}{dt}\right|_{t=0}\widetilde{S_{j}}[\omega_{t},\omega]=i\int_{M}j\operatorname{Td}_{j}(\nabla X,R,\cdot\cdot\cdot,R)\frac{\omega^{n+1-j}}{(n+1-j)!}+\operatorname{Td}_{j}(R,\cdot\cdot\cdot,R)\frac{\theta_{X}\omega^{n-j}}{(n-j)!}.$ (4.14)
On the other hand, we have
by dimensional reason. Comparing (4.14) and (4), we conclude
$\left.-\frac{d}{dt}\right|_{t=0}\widetilde{S_{j}}[\omega_{t},\omega]=\frac{i}{(n+1-j)!}\int_{M}\operatorname{Td}_{j}(R+\nabla X)(\omega+\theta_{X})^{n+1-j},$ (4.16)
which completes the proof. In the case of non-purely imaginary $\theta_{X}$, the proof is nearly identical to the one given above. ∎
Remark 4.
Note that $\theta_{X}$ is determined only up to an additive constant, but (4.7) behaves correctly under the addition of a constant to $\theta_{X}$. Using $S_{j}$ instead of $\widetilde{S_{j}}$ in the proof, a similar formula can be obtained which does not depend on the normalization of $\theta_{X}$ (for $j>0$).
For simplicity, we assume $\theta_{X}$ to be purely imaginary from now on. Recall that the Mabuchi functional integrates the Futaki invariant, which does not depend on Kähler metrics. Thus, (4.8) suggests the following proposition. Although this is a special case of [Fut04, Theorem 1.1] (by (4.7)), we include the proof for the reader’s convenience; see also [Tia00b, Remark 3.4].
Proposition 4.2.
Using the notation as above, for $j\geq 0$, $\int_{M}\theta_{X}\left(a_{j}(\omega)-\Delta a_{j-1}(\omega)\right)\frac{\omega^{n}}{n!}$ is independent of the choice of the Kähler metric in $\operatorname{c}_{1}(L)$, under the normalization of $\theta_{X}$ by $\int_{M}\theta_{X}\omega^{n}=0$.
Proof.
Let $\omega_{t}=\omega+i\partial\bar{\partial}\varphi_{t}$ be an arbitrary one-parameter family of Kähler forms in $[\omega]$. Define
$\widetilde{F_{j}}(\omega_{t},X):=\int_{M}\operatorname{Td}_{j}(R+\nabla X)(\omega+\theta_{X})^{n+1-j}.$ (4.17)
We will show that $\frac{d}{dt}\widetilde{F_{j}}(\omega_{t},X)=0$. Let $\theta_{t}$ be a potential function of $X$ with respect to $\omega_{t}$, that is, $\iota_{X}\omega_{t}=-\bar{\partial}\theta_{t}$. Setting $\alpha_{t}:=-i\partial\dot{\varphi_{t}}$, we have $\dot{\omega_{t}}=i\partial\bar{\partial}\dot{\varphi_{t}}=\bar{\partial}\alpha_{t}$ and
¯$\bar{\partial}\dot{\theta_{t}}=-\iota_{X}\dot{\omega_{t}}=-\iota_{X}\bar{\partial}\alpha_{t}=\bar{\partial}\iota_{X}\alpha_{t},$ (4.18)
where we used the fact that $X\in\mathfrak{h}(M)$. This implies $\dot{\theta_{t}}=\iota_{X}\alpha_{t}$ up to an additive constant. Also, by torsion freeness of the Levi–Civita connection we have
$\dot{\left(\nabla X\right)}=\dot{X^{q}_{p}}=\dot{\Gamma^{q}_{pr}}X^{r}=\iota_{X}\dot{\Gamma_{t}},$ (4.19)
@@
$\qquad\qquad\qquad\quad+(n+1-j)\operatorname{Td}_{j}(\nabla X+R)(\dot{\theta}+\dot{\omega})(\omega+\theta)^{n-j}$ (4.20)
@@
$\qquad+(n+1-j)\operatorname{Td}_{j}(\nabla X+R)(\iota_{X}\alpha+\bar{\partial}\alpha)(\omega+\theta)^{n-j}$ (4.21)
@@
where to get the last identity we used ¯$\bar{\partial}R=\bar{\partial}\omega=0$ and integration by parts.
@@
$\int_{M}\iota_{X}\left[j\operatorname{Td}_{j}(\dot{\Gamma},\nabla X+R,\cdots)(\omega+\theta)^{n+1-j}+(n+1-j)\operatorname{Td}_{j}(\nabla X+R)\alpha(\omega+\theta)^{n-j}\right]$ (4.25)
@@
Since ¯$\bar{\partial}\operatorname{Td}_{j}(\nabla X+R)=j\operatorname{Td}_{j}(\bar{\partial}\nabla X,\nabla X+R,\cdots)$, (4.22) and (4.26) are equal.
@@
$\frac{d}{dt}\widetilde{F_{j}}(\omega_{t},X)=\int_{M}\iota_{X}\eta=0$ (4.27)
by dimensional reason. ∎

### v5

@@ 3. Asymptotic expansion of the DPP partition function
$\tilde{S}_{\phi}[\varphi_{1},\varphi_{0}]:=\int_{M}BC(\phi;\omega_{1},\omega_{0})\frac{\omega_{0}^{n+1-j}}{(n+1-j)!}+\phi(R_{1})\frac{-i}{(n+1-j)!}\sum^{n-j}_{s=0}(\varphi_{1}-\varphi_{0})\omega_{1}^{s}\wedge\omega_{0}^{n-j-s}$
satisfies cocycle identity without polarization assumption on $\omega]$. Functionals of similar type were considered in [donaldson1985anti, donaldson1987infinite, tian1999bott]. Note that $\tilde{S}_{\phi}$ can be easily modified to become a functional on the space of Kähler forms, as we now demonstrate.
@@
This can be easily verified from either (3.3) or (3). We have an alternative expression for the asymptotics (3.6) where each of the coefficients is given by a geometric functional in the sense that it only depends on Kähler forms. That is, we prove Theorem 1. Let $V:=\int_{M}\omega^{n}/n!$ be a volume of $(M,\omega)$.
Proof of Theorem 1.
@@
$(2\pi)^{n}\log\frac{Z_{k}[\varphi]}{Z_{k}[0]}=kd_{k}(2\pi)^{n}S_{0}[\varphi,0]+k^{n}S_{1}[\varphi,0]+k^{n-1}S_{2}[\varphi,0]+\cdots,$
@@
$S_{0}[\varphi,0]:=\frac{1}{V}\widetilde{S_{0}}[\varphi,0]$
@@
$S_{j}[\varphi+c,0]=\widetilde{S_{j}}[\varphi+c,0]-\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L)\times\widetilde{S_{0}}[\varphi+c,0]$
@@
$\qquad-\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L)\times\widetilde{S_{0}}[\varphi,0]+c\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L)$
$=S_{j}[\varphi,0]$
for $j>0$. Thus, the functionals $S_{j}[\cdot,0],\,j>0$ on $\mathcal{K}_{\omega}$ descend to the functionals on Kähler metrics $\mathcal{K}_{0}=\mathcal{K}_{\omega}/\mathbb{R}$. Cocyle identity and exact cocycle property for $j>n+1$ follow from corresponding properties of $\widetilde{S_{j}}$, and (2) follows from (3.9). ∎
From (2.3) and (1.2), one can see that $S_{0}$ is (minus of) the Aubin-Yau functional $I$, and $S_{1}$ is the Mabuchi functional $\mathcal{M}$ in Kähler geometry. It can also be directly checked by comparing the explicit formula for $I,\mathcal{M}$ [tian2012canonical] with (3) and (3.9). For $j=2$, since $\beta_{1}$ only depends on $\omega]$, $S_{2}$ is explicitly given by
@@
$+\frac{i}{V}\left(\int_{M}\operatorname{Td}_{2}(T^{1,0}M)\frac{\omega^{n-2}}{(n-2)!}\right)\times\int_{M}\frac{-i}{(n+1)!}\sum^{n}_{s=0}\varphi\omega_{\varphi}^{s}\wedge\omega^{n-s}.$
@@
$\delta_{\varphi}S_{2}[\omega_{\varphi},\omega]=\frac{1}{6}\int_{M}\delta\varphi\Delta_{\varphi}R_{\varphi}\,\omega_{\varphi},$
which is the variation of the Liouville action restricted to fixed area metrics, written in terms of Kähler gauge [klevtsov2014random, (5.10)]. Therefore, $S_{2}$ extends the Liouville action to all dimensions. We will compute its second variation in Appendix A.
@@
In [klevtsov2014random, (6.6)], Klevtsov explicitly computed the asysmptotics of $\log\frac{Z_{k}[\varphi]}{Z_{k}[0]}$ up to $j\leq 4$ in dimension $n=1$. Based on that, it was conjectured that for the general dimension $n$, the first $n+2$ coefficients are nontrivial action functionals (satisfying cocycle identity) and terms with $j>n+1$ are exact cocycles. Thus, Theorem 1 confirms this conjecture. It was also asked what is the relationship between $S_{2}$ and known functionals in Kähler geometry, such as Chen-Tian functional $E_{1}$ [chen2002ricci]. In fact, from (3.10), one can see that $S_{2}$ is a linear combination of the Chen-Tian functional $E_{1}$ and the Bando-Mabuchi functional $\mathcal{M}_{2}$ [bando1986some]. By the results of [bando1986some, chen2002ricci], one can observe that $S_{2}$ integrates a Futaki-type holomorphic invariant. We will generalize this for all $j>0$ in the next section.
Note that (3.3) and (3) make sense as formal expressions for a nonpolarized Kähler class $\omega]$. We expect that they are equal in general, without a polarization assumption. In particular, it would imply that the 1-forms defined on $\mathcal{K}_{\omega}$ by
@@
for $\psi\in T_{\varphi}\mathcal{K}_{\omega}=C^{\infty}(M,\mathbb{R})$, are closed. It is known that this holds for $j=0,1$ by [mabuchi1986k], and for $j=2$ by [bando1986some, chen2002ricci]. We will verify these claims for $j=2$ in Appendix A by direct computation.
@@ 4. Holomorphic invariants from Bergman kernel asymptotics and critical points
In this section, we show that there is an obstruction to the existence of the critical points for each $S_{j}$, given by a holomorphic invariant introduced by Futaki [futaki2004asymptotic]. Proof of Theorem 2 is divided into two part, Theorem 4.1 and Theorem 4.2. Denote by $\mathfrak{h}(M)$ the Lie algebra of holomorphic vector fields on $(M,\omega)$. For $X\in\mathfrak{h}(M)$, by Hodge theory, there is a unique harmonic $(0,1)$-form $\tau$ and function $\theta_{X}$ (called holomorphy potential) satisfying
@@
Then $\theta_{X}$ is defined up to an additive constant, and $\tau$ will be assumed to be zero without loss of generality for our purposes.
If $\theta_{X}$ is real valued, ${\rm Re\,}X$ is a Hamiltonian vector field and $\frac{1}{2}\theta_{X}$ is a Hamiltonian function with respect to $\omega$, since
$\iota_{{\rm Re\,}X}\omega=-\frac{1}{2}(\bar{\partial}\theta_{X}+\partial\overline{\theta_{X}})=-\frac{1}{2}d\theta_{X}.$
That is, ${\rm Re\,}X$ lies in a Lie algebra of Hamiltonian symplectomorphisms $\mathcal{G}$ of $(M,\omega)$.
Consider $\nabla X=X^{q}_{p}\frac{\partial}{\partial z^{q}}\otimes dz^{p}$ as an $\operatorname{End}(T^{1,0}M)$-valued 0-form on $M$, where
$X^{q}_{p}=\frac{\partial X^{q}}{\partial z^{p}}+\Gamma^{q}_{pr}X^{r}.$
Denote by $\flat,\sharp$ the lowering and raising index operations by Kähler metric. For example,
$(\overline{\nabla X})^{\flat,\sharp}=g_{p\bar{r}}\overline{X^{r}_{l}}g^{q\bar{l}}\frac{\partial}{\partial z^{q}}\otimes dz^{p}.$
Let $R$ be the curvature 2-form of $\omega$. We introduce the following notation:
$\overline{R}:=(R)^{\flat,\sharp}=R^{\bar{q}}{}_{\bar{p}r\bar{l}}dz^{r}\wedge d\bar{z}^{l},$
which is $\operatorname{End}(T^{0,1}M)$-valued $2$-form on $M$.
First we identify the holomorphic invariants we will be dealing with. We begin with the lemma.
@@
We have
$\iota_{X}R=-\bar{\partial}\nabla X$ (4.2)
and
$\iota_{\overline{X}}\overline{R}=\partial\overline{\nabla X}.$ (4.3)
@@
$\bar{\partial}\nabla X$ $=\bar{\partial}\left(X^{q}_{p}\right)=\frac{\partial X^{q}_{p}}{\partial\bar{z}^{l}}d\bar{z}^{l}=\left(\frac{\partial^{2}X^{q}}{\partial z^{p}\partial\bar{z}^{l}}+\frac{\partial\Gamma^{q}_{pr}}{\partial\bar{z}^{l}}X^{r}+\Gamma^{q}_{pr}\frac{\partial X^{r}}{\partial\bar{z}^{l}}\right)d\bar{z}^{l}$
$=\frac{\partial\Gamma^{q}_{pr}}{\partial\bar{z}^{l}}X^{r}d\bar{z}^{l}=-R_{p}{}^{q}{}_{r\bar{l}}X^{r}d\bar{z}^{l}=-\iota_{X}R.$
Similarly,
$\partial\overline{\nabla X}=\partial\left(\overline{X^{q}_{p}}\right)=\frac{\partial\overline{\Gamma^{q}_{pl}}}{\partial z^{r}}\overline{X^{l}}dz^{r}=\overline{-R_{p}{}^{q}{}_{l\bar{r}}X^{l}}dz^{r}=-R^{\bar{q}}{}_{\bar{p}r\bar{l}}\overline{X^{l}}dz^{r}=\iota_{\overline{X}}\overline{R}$
where we used the identity
$\overline{R_{i\bar{j}k\bar{l}}}=R_{j\bar{i}l\bar{k}}.$
Using the lemma, we can prove the invariance of the following invariants.
@@
Under the same notations as above, for $j\geq 0$,
$\widetilde{F_{j}}(\omega,X):=\int_{M}\operatorname{Td}_{j}(R+\nabla X)(\omega+\theta_{X})^{n+1-j}$ (4.4)
and
$\widetilde{F_{j}}(\omega,\overline{X}):=\int_{M}\operatorname{Td}_{j}(R+(\overline{\nabla X})^{\flat,\sharp})(\omega-\overline{\theta_{X}})^{n+1-j}$ (4.5)
are independent of the choice of the Kähler metric in $\operatorname{c}_{1}(L)$, under the normalization of $\theta_{X}$ by $\int_{M}\theta_{X}\omega^{n}=0$.
@@
We first prove the claim for (4.4). The proof for (4.5) is exactly the same, as we will explain.
Let $\omega_{t}=\omega+i\partial\bar{\partial}\varphi_{t}$ be an arbitrary one-parameter family of Kähler forms in $\omega]$ with $\omega_{0}=\omega$. We will show that $\frac{d}{dt}\widetilde{F_{j}}(\omega_{t},X)=0$. Let $\theta_{t}$ be a holomorphy potential of $X$ with respect to $\omega_{t}$, that is, $\iota_{X}\omega_{t}=-\bar{\partial}\theta_{t}$. Setting $\alpha_{t}:=-i\partial\dot{\varphi_{t}}$, we have $\dot{\omega_{t}}=i\partial\bar{\partial}\dot{\varphi_{t}}=\bar{\partial}\alpha_{t}$. Since
$\iota_{X}\omega_{t}=\iota_{X}\omega+i\iota_{X}\partial\bar{\partial}\varphi_{t}=-\bar{\partial}\theta_{0}+i\bar{\partial}\iota_{X}(\partial\varphi_{t}),$
we have $\theta_{t}=\theta_{0}-i\iota_{X}\partial\varphi_{t}$ and thus $\dot{\theta_{t}}=\iota_{X}\alpha_{t}$ up to an additive constant. We used the fact that $X\in\mathfrak{h}(M)$. Also, by torsion freeness of the Levi–Civita connection we have
$\dot{\left(\nabla X\right)}=\dot{X^{q}_{p}}=\dot{\Gamma^{q}_{pr}}X^{r}=\iota_{X}\dot{\Gamma_{t}},$
@@
$\qquad\qquad\qquad\quad+(n+1-j)\operatorname{Td}_{j}(\nabla X+R)(\dot{\theta}+\dot{\omega})(\omega+\theta)^{n-j}$
@@
$\qquad+(n+1-j)\operatorname{Td}_{j}(\nabla X+R)(\iota_{X}\alpha+\bar{\partial}\alpha)(\omega+\theta)^{n-j}$
@@
where to get the last identity we used $\bar{\partial}R=\bar{\partial}\omega=0$ and integration by parts.
@@
$\int_{M}\iota_{X}\left[j\operatorname{Td}_{j}(\dot{\Gamma},\nabla X+R,\cdots)(\omega+\theta)^{n+1-j}+(n+1-j)\operatorname{Td}_{j}(\nabla X+R)\alpha(\omega+\theta)^{n-j}\right]$
@@
Since $\bar{\partial}\operatorname{Td}_{j}(\nabla X+R)=j\operatorname{Td}_{j}(\bar{\partial}\nabla X,\nabla X+R,\cdots)$, (4) and (4) are equal.
@@
$\frac{d}{dt}\widetilde{F_{j}}(\omega_{t},X)=\int_{M}\iota_{X}\eta=0$
by dimensional reason.
We now turn to the proof for (4.5). Note that
$\operatorname{Td}_{j}(R+(\overline{\nabla X})^{\flat,\sharp})=\operatorname{Td}_{j}(\overline{R}+\overline{\nabla X}),$
where on the right hand side trace is taken as $\operatorname{End}(T^{0,1}M)$. Then as the proof for (4.4) shows, we only need to check the following identities:
$\dot{\overline{R}}=-\partial\dot{\overline{\Gamma}};\quad\dot{\overline{\nabla X}}=\iota_{\overline{X}}\dot{\overline{\Gamma}};\quad\dot{\omega}=\partial\overline{\alpha};\quad\dot{\overline{\theta}}=\iota_{\overline{X}}\overline{\alpha},$
and
$\iota_{\overline{X}}(\omega-\overline{\theta})=-\partial\overline{\theta};\quad\iota_{\overline{X}}(\overline{R}+\overline{\nabla X})=\partial\overline{\nabla X}.$
For example,
$\dot{\overline{R}}=\dot{R}^{\bar{q}}{}_{\bar{p}r\bar{l}}dz^{r}\wedge d\bar{z}^{l}=\dot{\overline{R_{p}{}^{q}{}_{l\overline{r}}}}dz^{r}\wedge d\bar{z}^{l}=\overline{-\partial_{\bar{r}}\dot{\Gamma^{q}_{pl}}}dz^{r}\wedge d\bar{z}^{l}=-\partial_{r}\dot{\overline{\Gamma^{q}_{pl}}}dz^{r}\wedge d\bar{z}^{l}=-\partial\dot{\overline{\Gamma}}.$
The proof then goes without changes, and one can show that the time derivative is equal to
$\int_{M}\iota_{\overline{X}}\left[j\operatorname{Td}_{j}(\dot{\overline{\Gamma}},\overline{\nabla X}+\overline{R},\cdots)(\omega-\overline{\theta})^{n+1-j}-(n+1-j)\operatorname{Td}_{j}(\overline{\nabla X}+\overline{R})\overline{\alpha}(\omega-\overline{\theta})^{n-j}\right],$
which is zero. Note the minus sign in front of the second term. ∎